# Cross-Encoder Rediscovers a Semantic Variant of BM25

### v1

@@ 4.1. Relevance Scoring Heads
Figure 4. Example of the attention pattern from [CLS] to query tokens in the Relevance Scoring Heads, illustrating how these heads process soft-TF for specific query tokens based on their IDF values.
Component Behavior. Upon closer inspection of the attention patterns of these heads, we observe that the [CLS] token selectively attends to query tokens and retrieves soft-TF based on each query token’s IDF value (Figure 4). Specifically, head 10.1 focuses predominantly on the highest-idf query token: the more uncommon a query token (i.e., the higher its IDF), the more the [CLS] token attends to it to capture its soft-TF information. Conversely, head 10.10 primarily attends to low-IDF tokens, while heads 10.7 and 10.4 focus on middle-to-high-IDF tokens.
To quantify this observation, we first compute the Pearson correlation between each head’s attention pattern and the query tokens’ IDF values. The results show strong correlations for heads 10.1 ($r=0.878$), 10.4 ($r=0.780$), and 10.7 ($r=0.798$), with $p<0.001$ for all, establishing a statistically significant relationship (threshold: $p<0.05$) between the attention values from [CLS] and the query tokens’ IDF values for 10.1, 10.4 and 10.7. However, Head 10.10 exhibits a weaker correlation ($r=0.206$). Then, we calculate the average IDF value weighted by each head’s attention distribution across query tokens to further validate this trend. Head 10.1 has the highest weighted average IDF (2.704), followed by heads 10.7 (1.983), 10.4 (2.171), and finally head 10.10 (1.457). This ranking quantitatively supports our observation and explains the weaker correlation between 10.10’s attention value with IDF values due to its negative association with them. Thus, each head covers different parts of the query based on IDF ranges, and together they send soft-TF information for all the query tokens to [CLS] together using a computation consisting of both soft-TF and IDF.
BM25-based Scoring Hypothesis. Considering that the mentioned computation resembles BM25, we hypothesize that the Relevance Scoring Heads perform a BM25-like summation of soft-TF values weighted by IDF of the query tokens. The model might attribute importance to the output of 10.1, 10.4 & 10.7, 10.10 from highest to lowest (e.g. controlled by different output projection matrix values) such that if a query token’s soft-TF is processed by 10.1, its soft-TF contributes the most to the relevance score. In this way, the query token’s soft-TF contribution to the relevance score is proportional to its IDF value. We verify this hypothesis in §5.
Figure 5. Example attention pattern of Matching Head 4.9: tokens attend most strongly to duplicates but also mildly to similar tokens.
@@
Important Upstream Heads. Continuing our “backward pass,” we path patch to the relevance scoring heads to identify upstream components transmitting soft-TF information. This patching results for TFC1 and STMC1 show high correlation of 1.0 ( p ¡0.001) and reveals a shared set of attention heads, divided into two functional groups: (1) Contextual Query Representation Heads4.2) in middle layers, which reorganize soft-TF information among query tokens to construct a contextualized query representation. (2) Matching Heads4.3) in early layers, which compute Matching Scores (soft-TF weighted by term saturation and document length effects).
4.2. Contextual Query Representation Heads
@@
The first group of heads that send information to Relevance Scoring Heads are termed contextual query representation heads: 8.10 and 9.11. These heads aggregate soft-TF information of the higher-IDF query tokens and distribute this information among all query tokens, strengthening their representations for the Relevance Scoring Heads’ final computation.
@@
At the Relevance Scoring Heads, the [CLS] token retrieves soft-TF information from all query tokens. Thus, we analyze the attention patterns among query tokens in these two intermediary heads to understand how they modify the query token representation in the residual stream. The attention patterns reveal a distinct pattern: all query tokens in these heads consistently focus on one or two higher-IDF query tokens, with strong correlations to IDF values (9.11: $r=0.829$, 8.10: $r=0.781$ ) with both $p<0.001$. The evidence suggests that 8.10 and 9.11 act as Contextual Query Representation Heads: they distribute the soft-TF information of higher-IDF tokens across all query tokens. This ensures that when [CLS] token retrieves soft-TF information at Relevance Scoring Heads, each query token retains a stronger representation of the higher-IDF tokens’ soft-TF values.
@@
Important Upstream Components. Path patching to 8.10 and 9.11 confirms that these heads also receive soft-TF information from the Matching Heads.
4.3. Matching Heads
Matching Heads (0.8, 1.7, 2.1, 3.1, 4.9, 5.7, 5.9, 6.3, 6.5, 7.9, 8.0, 8.1, 8.8) compute and transmit a Matching Score, combining soft-TF, term saturation, and document length signals.
Component Behavior. Patching results reveal that Matching Heads transmit soft-TF information to the Relevance Scoring Heads and Contextual Query Representation Heads. A key question, however, is whether these heads merely propagate soft-TF signals from earlier components or compute the soft-TF themselves. Analyzing their attention patterns indicates that they actively attend to semantic matches and compute the soft-TF. Figure 5 illustrates this behavior for head 1.7: all query and documents tokens are strongly attending duplicates and mildly attending to semantically-similar terms.
Qualitative examination of the 13 heads with positive patching importance reveals attention patterns similar to those in Figure 5. We quantitatively check whether the query tokens are in fact matching semantically by measuring the correlation between the attention values (from any token to any document token) and the tokens’ semantic similarity, approximated using cross-encoder’s embeddings. The results show that Matching Heads exhibit a significantly higher average Pearson correlation (corr = 0.500, p¡0.001) compared to other heads (corr = 0.132), confirming that these heads function as Matching Heads: at these heads, query tokens actively attend to semantically similar terms, with the attention value serving as an indicator of semantic similarity. Following this intuition, the sum of attention values from a query token to all document tokens can approximate the soft-TF of a query term.
To validate the proposed function of the Matching Heads, we mean-ablate these Matching Heads and observe that the TF perturbed samples’ average value drops from 5.146 to -4.394, and the STMC perturbed samples’ average value drops from -1.998 to -4.611. The drastic drop occurs because ablation deactivates all semantic-matching functions for all query terms.
Figure 6. For the majority of Matching Heads, after the initial occurrence of $X_{\text{selected}}$, which causes a sharp increase in attention, subsequent occurrences result in only minimal incremental increases. This aligns with TFC2, which states that additional term occurrences yield smaller improvements.
Figure 7. Attention trends for selected terms are shown, with unselected terms following a similar pattern: longer documents generally increase the average attention assigned to all document tokens, regardless of whether they match a query token.
Matching Score Contains Term Saturation and Document Length Signals. The average Soft-TF Correlation Score of Matching Heads (0.500) indicates that their attention values from query tokens to document tokens encode not only semantic similarity but also additional signals. To examine whether these attention values capture term saturation or document length effects, we analyze their behavior under using two diagnostic datasets: (1) TFC2 (increasing occurrences of a query term), and (2) LNC1 dataset (increasing number of irrelevant sentences to simulate the isolated effect of longer document length).
As the occurrences of $x_{\text{selected}}$ increase, its average attention value grows (Figure 6). On the other hand, the attention for $x_{\text{others}}$ remains relatively constant, which aligns with our soft-TF understanding: only document tokens that are semantically similar to a query token get nonzero attention values, proportional to the extent and frequency of semantic similarity. Furthermore, for most heads, the TFC2 trend is preserved in the attention values. For LNC1, injecting more irrelevant sentences leads to an increase in the attention value of all query tokens, regardless of whether they match specific query terms (Figure 7). Notably, this increase in attention occurs even as the overall relevance score decreases (as expected from LNC1), suggesting that attention values encode mixed and composite signals of soft-TF, term saturation, and document length, which the model has learned to disentangle to produce appropriate relevance score changes. The varying degrees of influence on the heads’ attention values shown in Figure 6 and 7 suggest that some Matching Heads play a more significant role in regulating these effects, and they may collectively approximate the ideal term effects. Thus, we define this complex attention value as the Matching Score to reflect that it encapsulates soft-TF, term saturation, and document length signals, three important signals of BM25.
Important Upstream Components. Due to time constraint, we further path patch to the three most important Matching Heads (7.9, 4.9, 2.1) to track how soft-TF flows into them. We find that the embedding exhibits the most direct influence, while previous attention and mlp layers exhibit very little influence. This completes our tracing of soft-TF path, confirming that the Matching Heads are the original “generator” of soft-TF as query token attends to soft matches to compute the Matching Score.
4.4. IDF in the Embedding Matrix
In addition to soft-TF, IDF is a critical signal for Relevance Scoring Heads. Here, we use SVD and low-rank intervention to track where IDF is stored or computed.
IDF exists in the first low-rank component of the embedding matrix. Previously, Choi et al.  (Choi et al., 2022) provide correlational evidence of IDF information existing in the embedding matrix. We decompose the embedding space into orthogonal components using Singular Value Decomposition (SVD). Recall that the embedding matrix $W_{E}$ is structured such that each row corresponds to a word’s representation in the vocabulary. The columns of $W_{E}$ represent different dimensions of this vector space, capturing various latent features of the words.
SVD allows us to decompose a complex matrix into a sum of rank-1 matrices, ordered by their contribution to the overall structure of the matrix. A rank-1 matrix is the outer product of a column and row vector, capturing a single direction in the data. This decomposition enables us to represent $W_{E}$ in orthogonal directions, allowing us to analyze the dominant directions in an easily interpretable manner due to their rank-1 structure. Mathematically, the SVD of $W_{E}$ is expressed as:
@@ 4.4. IDF in the Embedding Matrix
Here, $\sigma_{i}$ are the singular values, which quantify the importance or strength of each rank-1 component; $u_{i}$ and $v_{i}$ left and right singular vectors, respectively, forming orthonormal bases that capture patterns in the row (token embeddings) and column spaces (feature dimensions); and $r$ is the rank of $W_{E}$, representing the number of non-zero singular values and independent directions in the matrix. By focusing on the largest singular values ($\sigma_{i}$) and their corresponding singular vectors ($u_{i},v_{i}$), we can study whether IDF is stored in dominant components of the Embedding.
Notably, we find that the rank-1 vector with the highest singular value, we call $U_{0}$, has a large negative pearson correlation ($-71.36\%$) with the IDF value from the MS-MARCO train set (Nguyen et al., 2016): higher values in this low-rank component for a given vocabulary token correspond to lower IDF values for that token. This suggests that IDF information is encoded prominently within the most significant low-rank component of the embedding matrix.
To validate this correlation, we perform an intervention in the next section to demonstrate that the IDF values from $U_{0}$ have a causal effect on downstream components (i.e., relevance scoring heads) and thus, the overall relevance computation.
@@
Figure 8. $U_{0}$ editing example. Since $U_{0}$ is negatively correlated with IDF, increasing token 1’s (tok1) IDF, representing its importance in relevance computation, requires scaling its value in $U_{0}$ negatively. We transpose $U_{0}$ for visualization.
@@
IDF intervention shows causal evidence that the model uses IDF in $U_{0}$ for relevance computation. Given previous understanding on $U_{0}$, we can interpret $U_{0}$ as a 1-D IDF dictionary, where $\text{idx}(q_{i})$ corresponds to the vocabulary index of $q_{i}$, mapping to its respective IDF value. Modifying the value at $\text{idx}(q_{i})$ in $U_{0}$ allows us to adjust the importance of the Matching Score for $q_{i}$ (Figure 8).
@@
If the model uses the IDF values encoded in $U_{0}$, then modifying these values should result in corresponding changes to the ranking score. Specifically, according to the BM25-based Scoring Hypothesis in §4.1, there is a linear relationship between the IDF of a query token $q_{i}$ and the relevance score: increasing the IDF of $q_{i}$ should increase the relevance score. We have shown that $U_{0}$ is rank-1 and negatively correlated with IDF. If the model indeed uses the IDF values encoded in $U_{0}$, we can increase $\text{IDF}(q_{i})$ by decreasing qi’s value in U0, which would directly increase the relevance score. The converse should hold true for decreasing IDF.
@@
One way to observe whether this is happening is to look at very controlled, minimal examples that we curate specifically to fully observe the effect. Using our base dataset, we create two documents, doc1 and doc2, containing distinct repeated tokens: first token (tok1) and second token (tok2), respectively (e.g., query: “computer science department number”, doc1: “computer computer computer”, doc2: “science science science”). Then, we can edit the IDF of tok1 and evaluate the IDF editing effect on the doc1 and doc2 relevance scores.
@@
Figure 9. Top: Changes in scores for doc1 (repeated tok1) and doc2 (repeated tok2) when increasing tok1’s IDF at different scales, averaged across all samples. Larger increases in tok1’s IDF result in larger score increase for doc1. Bottom: Changes in scores for doc1 and doc2 when decreasing tok1’s IDF shows the inverse effect. Both increase and decrease has ranges where the monotonic trend reverses a little which suggests that there are optimal range for editing both ways.
Figure 9 offers support for the direct causal relationship: across most scaling factors, larger increase (or decrease) in tok1’s IDF result in larger score increase (or decrease, respectively) for doc1. The largely monotonic trend not only offers causal evidence for the model’s use of $U_{0}$ to acquire IDF values but also shows preliminary evidence supporting that the model computes a sum of soft-TF weighted by IDF values (§4.1).
Furthermore, the discovery of the low-rank nature of the IDF signals enables targeted interventions. Our intervention experiment demonstrates that modifying IDF values allows for precise control over the importance of specific terms within the cross-encoder.
5. Validation of BM25-like Computation
After we established where all the BM25-like components were initially computed or stored and how they were passed to Relevance Scoring Heads, we now assess whether the Relevance Scoring Heads perform a BM25-style computation as hypothesized in §4.1. To do this, we first formalize the hypothesized function of the heads and the information flowing into them as a BM25-style linear function. Next, we evaluate the linear model’s ability to reconstruct the cross-encoder scores by examining how well the hypothesized linear model fits the data.
5.1. Formalizing the Hypothesized Function
Specifically, as proposed in the BM-based Scoring Hypothesis, the relevance scoring heads appear to compute a summation of Matching Scores weighted by IDF. The Matching Score incorporates soft-TF, along with term saturation and document length signals, all of which are components of the BM25 function. If this hypothesis holds—that these components interact in a BM25-like manner—then the semantic scoring circuit can be expressed as a linear function:
$\sum_{i=1}^{N}\text{linear\_combo}\big{(}-U_{0}(q_{i}),\text{MS}_{\text{total}}(q_{i},d_{j}),[-U_{0}(q_{i})\cdot\text{MS}_{\text{total}}(q_{i},d_{j})]\big{)}$

### v2

@@ 4.1 Relevance Scoring Heads
Figure 3: Example of the attention pattern from [CLS] to query tokens in the Relevance Scoring Heads, illustrating how these heads process soft-TF for specific query tokens based on their IDF values.
Component Behavior. Qualitative analysis reveals that the [CLS] token selectively attends to query tokens, indicating the movement of soft-TF signals into its representation to prepare for the final relevance score in the model’s classification head. We find the heads distribute attention to different parts of the query. Specifically, 10.1 focuses on high-IDF query terms, 10.10 on low-IDF, and 10.7 and 10.4 on mid-to-high-IDF (Figure 3). We verify this distributive behavior by calculating the Pearson correlation between each head’s attention distribution and the IDF values of query tokens and find a moderately high average (r=0.67, p < 0.01).
Comparison to BM25. Because these heads prepare the [CLS] token for the final scoring in the classification head, we call them Relevance Scoring Heads. They combine IDF and soft-TF signals per query term, resembling BM25’s term weighting mechanism. Thus, we hypothesize that the model combines the outputs of the Relevance Scoring Heads in a similar manner and verify this hypothesis in §5.
@@
Important Upstream Components. We path patch to the Relevance Scoring Heads’ value vectors to identify which upstream components transmit soft-TF signals. On both TFC1 and STMC1, patching effects show high average correlation of 0.83 (p <0.001), revealing a shared set of attention heads, divided into two groups: (1) Query Contextualization Heads (§4.2) which redistribute soft-TF among query tokens and (2) Matching Heads (§4.3) which detect exact and semantic query-document matches (soft-TF) (see Appendix C).
4.2 Query Contextualization Heads
@@
Query Contextualization Heads (8.10 and 9.11) aggregate soft-TF signals of higher-IDF query tokens and distribute them across all query tokens, strengthening their representations for the Relevance Scoring Heads’ final computation.
@@
Component Behavior. At the Relevance Scoring Heads, the [CLS] token retrieves soft-TF from all query tokens. Thus, we analyze the attention patterns among query tokens in these two intermediary heads to understand how they modify the query token representation in the residual stream. We find that all query tokens in these heads consistently focus on one or two higher-IDF query tokens, with strong correlations to IDF values (9.11: $r=0.829$, 8.10: $r=0.781$ ) with both $p<0.001$. This suggests that 8.10 and 9.11 redistribute the soft-TF signals of higher-IDF tokens to contextualize the entire query for the Relevance Scoring Heads.
Comparison to BM25. BM25 weights each query term independently, with no exchange of information between terms. In contrast, Query Contextualization Heads seem to learn to contextualize the entire query, redistributing soft-TF signals to amplify high-IDF tokens and dynamically reweight terms, before the final scoring. Further study of this learned contextualization is left for future work.
@@
Important Upstream Components. Path patching to 8.10 and 9.11 confirms that these heads receive soft-TF signals from the Matching Heads (Appendix C).
4.3 Matching Heads
Figure 4: Example attention pattern of Matching Head 4.9: tokens attend most strongly to duplicates but also mildly to similar tokens.
Alongside Query Contextualization Heads, a separate set of heads, named Matching Heads (0.8, 1.7, 2.1, 3.1, 4.9, 5.7, 5.9, 6.3, 6.5, 7.9, 8.0, 8.1, 8.8), in the model’s early and middle layers detect exact and semantic token matches (soft-TF) between the query and document and pass these signals to the Relevance Scoring Heads.
Component Behavior. Qualitative inspection reveals that Matching Heads actively attend to both exact and semantically similar matches across the entire input sequence (i.e., query and document). As shown in Figure 4, head 1.7 exhibits this behavior clearly: both query and document tokens strongly attend to duplicated terms and, to a lesser extent, to semantically related terms.
We compute the Pearson correlation between attention scores and semantic similarity for each query-document token pair, with semantic similarity measured as the cosine similarity between token embedding vectors. Matching Heads on average exhibit a significantly stronger correlation (r = 0.500, p < 0.001) than other heads (r = 0.132). This evidence suggests that Matching Heads employ an attention mechanism that scales proportionally with semantic proximity: higher semantic similarity corresponds to higher attention scores. Since each attention weight quantifies the strength of token matching, the sum of attention values from a query token to all document tokens effectively approximates the soft-TF of a query term.
We validate the Matching Heads’ importance by mean-ablating them, and find that the average logits across both TFC1 and STMC1 perturbed samples decrease (TFC1: 5.146 to -4.394, STMC1: -1.998 to -4.611), indicating that these heads are critical for computing a semantic version of term matches.
Additional Relevance Signals. Since BM25’s TF term takes two additional signals, term saturation and document length, into account, we also investigate whether they affect Matching Heads. To isolate these effects, we use two diagnostic datasets: (1) TFC2, which increases repeated query term occurrences to test for term saturation, and (2) LNC1, which adds irrelevant sentences to simulate the isolated effect of increased document length (Table 1).
Because each dataset varies only in one factor, we track each Matching Head’s average attention across these controlled groups. Consistent with how BM25 controls for term saturation and document length effects, we observe two trends (see Appendix D): (1) in TFC2, attention to a term increases sharply after the initial occurrence, but plateaus with additional term repetitions, consistent with term saturation, and (2) in LNC1, attention decreases as irrelevant content increases, despite constant relevance, indicating sensitivity to document length. These findings suggest that Matching Heads integrate soft-TF with saturation and document-length signals rather than operating in isolation. We define this composite signal as the Matching Score, reflecting its incorporation of soft term frequency, term saturation, and document length effects, three core signals of BM25.
Comparison to BM25. Matching Heads implement a semantic variant of the TF component in BM25 by identifying and weighting query-document token matches. Our TFC2 and LNC1 experiments show that they also go beyond simple match counting: their attention outputs are modulated in ways that mirror BM25’s additional relevance adjustments for term saturation and document length normalization. Thus, Matching Heads jointly capture all three core components of BM25, TF, term saturation, and length normalization, while also incorporating semantic similarity.
Important Upstream Components. Path patching reveals that the embeddings are the primary input to Matching Heads (see Appendix C.2), confirming they generate soft-TF based on semantic similarity from the embeddings. This completes the backward tracing for the soft-TF circuit.
4.4 IDF in the Embedding Matrix
In addition to TF, IDF plays a critical role in BM25 by weighting each query token’s TF contribution, allowing the model to prioritize more uncommon terms. As shown in §4.1, IDF similarly emerges as a key signal for the Relevance Scoring Heads in our model. While Choi et al. (2022) provide correlational evidence for IDF in the embedding matrix, we use SVD and low-rank interventions to causally localize their functional role in the model.
SVD allows us to decompose a matrix into a sum of orthogonal rank-1 components, outer products of a column and row vector, ordered by their contribution to the overall matrix. Recall that the embedding matrix $W_{E}$ is structured such that each row corresponds to a word’s representation in the vocabulary, while the columns represent latent feature dimensions. By applying SVD to $W_{E}$, we can analyze the dominant directions to see what signals (e.g., IDF) drive the model’s behavior. Mathematically, the SVD of $W_{E}$ is expressed as:
@@ 4.4 IDF in the Embedding Matrix
Here, $\sigma_{i}$ are the singular values, which quantify the importance or strength of each rank-1 component; $u_{i}$ and $v_{i}$ left and right singular vectors, respectively, forming orthonormal bases that capture patterns in the row (token embeddings) and column spaces (feature dimensions); and $r$ is the rank of $W_{E}$, representing the number of non-zero singular values and independent directions in the matrix. By focusing on the largest singular values ($\sigma_{i}$) and their corresponding singular vectors ($u_{i},v_{i}$), we can study whether IDF is stored in dominant components of the embedding matrix.
We find that the top singular vector $U_{0}$ is highly correlated (r = $-71.36\%$) with MS-MARCO IDF values Nguyen et al. (2016), the model’s training dataset, indicating that IDF is encoded in the embedding matrix’s dominant low-rank component.
@@
Figure 5: $U_{0}$ editing example (transposed for visualization). Since $U_{0}$ is negatively correlated with IDF, increasing tok1’s IDF, representing its importance in relevance computation, requires negatively scaling its $U_{0}$ component.
Causal Experiment. To validate this correlation, we perform an intervention to demonstrate that the IDF values from $U_{0}$ have a causal effect on downstream components (i.e., relevance scoring heads) and thus, the overall relevance computation.
@@
Given previous understanding on $U_{0}$, we can interpret $U_{0}$ as a 1-D IDF dictionary, where $\text{idx}(q_{i})$ corresponds to the vocabulary index of $q_{i}$, mapping to its respective IDF value. Modifying the value at $\text{idx}(q_{i})$ in $U_{0}$ allows us to adjust the importance of the Matching Score for $q_{i}$ (Figure 5).
@@
If the model uses the IDF values encoded in $U_{0}$, then modifying these values should result in corresponding changes to the ranking score. Specifically, according to the BM25-based Scoring Hypothesis in §4.1, there is a linear relationship between the IDF of a query token $q_{i}$ and the relevance score: increasing the IDF of $q_{i}$ should increase the relevance score. We have shown that $U_{0}$ is rank-1 and negatively correlated with IDF. If the model indeed uses the IDF values encoded in $U_{0}$, we can increase $\text{IDF}(q_{i})$ by decreasing $q_{i}$’s value in $U_{0}$, which would directly increase the relevance score. The converse should hold true for decreasing IDF.
@@
To test this, we design an experiment with controlled, minimal examples. Given a query from our base dataset (e.g., “computer science department number”), we create two documents where doc1 repeats the first query token (e.g., “computer computer computer” and doc2 repeats the second query token (e.g., “science science science”). Then, we can edit the IDF of tok1 and measure the effect on the relevance scores of doc1 and doc2.
@@
Figure 6: Top: Changes in scores for doc1 (repeated tok1) and doc2 (repeated tok2) when increasing tok1’s IDF at different scales, averaged across all samples. Increasing tok1’s IDF raises doc1’s score more than doc2’s. Bottom: Decreasing tok1’s IDF produces the inverse effect, with slight nonmonotonic deviations suggesting optimal editing windows.
Figure 6 shows that scaling tok1’s IDF up or down causes a corresponding monotonic increase (or decrease) in doc1’s score, providing causal evidence that the model uses $U_{0}$ to encode IDF and suggests that it sums soft-TF by IDF values (§4.1).
Comparison to BM25. This completes the BM25 component set: IDF embeddings encode IDF, and Relevance Scoring Heads modulate the amount of soft-TF extracted by the [CLS] token based on that IDF. This is similar to BM25, where each query term’s TF is weighted by its IDF to prioritize informative terms. In the next section, we formally validate that Relevance Scoring Heads combine soft-TF and IDF in a BM25 manner to compute the final relevance score.
5 Validation of BM25-like Computation
To validate whether Relevance Scoring Heads perform a BM25-style computation as hypothesized in §4.1, we first formalize the hypothesized function of the heads and the information flowing into them as a BM25-style linear function. Next, we evaluate the linear model’s ability to reconstruct the cross-encoder scores by examining how well the hypothesized linear model fits the data.
5.1 Formalizing the Hypothesized Function
In §4.1, we hypothesize that Relevance Scoring Heads appear to compute a summation of Matching Scores weighted by IDF. The Matching Score incorporates soft-TF, along with term saturation and document length signals, all of which are components of the BM25 function. If this hypothesis holds, that these components interact in a BM25-like manner, then the semantic scoring circuit can be expressed as a linear function:
[…]
$\quad-\,U_{0}(q_{i})\,\cdot\,\text{MS}_{\mathrm{total}}(q_{i},d_{j})\Bigr{)}$